OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 30 - 6n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 4 practice quizzes? 20 3 5 6

OpenStudy (ash2326):

The no. of incorrect answers is given as \[30-6n\] n= no of practice quizzes he took It's given he took 4 quizzes so n=4 \[30-6 \times 4\] \[= 30-24=36\] So he got 36 wrong

OpenStudy (ash2326):

sorry What I'm doing:( 30-24=6 so he got 6 wrong

OpenStudy (anonymous):

IT'S OKAY I KNO WUD YHU WAS TRYN TO SAT